If P(E)=0.45, P(E or F)=0.50 and P(E and F)=0.20, find P(F)
Answer by jim_thompson5910(35256)

P(E or F) = P(E) + P(F) - P(E and F)

0.50 = 0.45 + P(F) - 0.20

0.50 = 0.25 + P(F)

0.50 - 0.25 = P(F)

0.25 = P(F)

P(F) = 0.25